OpenStudy (anonymous):

Find the complex solutions of the equation: 3x^2-5x+7=0

OpenStudy (luigi0210):

Use the quadratic formula

OpenStudy (anonymous):

what does it mean by complex solutions though

OpenStudy (anonymous):

Complex solutions are those that may be written in the form \(a+bi\) with \(a,b\) real and \(i=\sqrt{-1}\).

OpenStudy (anonymous):

Since it's a second degree polynomial, we know there will be two complex solutions (or only one with multiplicity 2).

OpenStudy (anonymous):

25 you mean not 24

OpenStudy (anonymous):

$$x=\frac{5\pm\sqrt{25-84}}6=\frac{5\pm\sqrt{-59}}6=\frac{5\pm i\sqrt{59}}6$$

OpenStudy (anonymous):

Are the two solutions 2.1 or -0.45?

OpenStudy (anonymous):

They are "complex solutions", not real numbers. Complex solutions implies that the solutions are complex numbers.

OpenStudy (anonymous):

so how would i go about finding the solutions then?